A 4.0 kg mass has a velocity of 10 m/s to the EAST. The 4.0 kg mass undergoes an acceleration of 4.0 m/s^2 to the WEST for 3.0 seconds. What is the velocity of the 4.0 kg mass at the end of the 3.0 sec interval?
A) 22m/s to the west
B) 2.0 m/s to the west
C) 0.0 m/s
D) 2.0 m/s to the east
E) 22 m/s to the east
Answer #1

v = u - at

v = 10 - 4*3

v = 10 - 12

v = -2 m/s

Velocity = 2.0 m/s to the west

the velocity of the 4.0 kg mass at the end of the 3.0 sec interval is 2 m/s to the west.